# Integrating Spatiotemporal Features in LSTM for Spatially Informed COVID-19 Hospitalization Forecasting

### v1

@@ 3.5 Model Training
We employed a walk-forward validation approach to train the model, utilizing all available data up to and including the forecast date, which is made accessible before the Hub submission deadline every week. Models for each forecast date are trained using the most recent 15 months of data, capturing the latest transmission patterns, hospitalization dynamics, and population immunity. We use spatial cross-validation for training our models, and the final testing is done purely on forecasting dates during periods occurring after the training set period. We configured our training strategy such that for each forecast date, it selects three states exhibiting the lowest, median, and highest hospitalization rates (a criterion that changes weekly, so different states occupy each role over time) to serve as our validation for early stopping. This spatial validation strategy ensures that the model is trained while consistently validated against a diverse range of epidemic dynamics- specifically, states at increasing, decreasing, and persisting stages of epidemic spread (Lever et al. 2016). This ensures that the model is not overfitting to increasing or decreasing trends. The remaining states, with varying hospitalization rate trends are used for training the model. This method allows us to implement early stopping to prevent overfitting and test the model’s generalizability across a representative cross-section of epidemic conditions. After training is completed, the model is tested on time periods following the training period (with no overlap), to ensure the model’s ability to forecast the future.
During training, the time-series SPH values, along with incident cases and hospitalization rates, are first normalized to the range [-1, 1] using the MinMax Scaling technique. These normalized features are then sequentially fed into the LSTM model, which processes the three input features over successive time steps. This integration helps to enhance the model’s ability to predict hospitalizations based on not only historical patterns of the same state, but also historical patterns of connected states, with the connectedness strength measured by social connectivity.
@@ 3.6 Model Evaluation
While SLSTM was employed during various phases of the pandemic to submit forecasts to the Hub, for the results presented in this paper, we utilized data from July 27, 2020, to March 7, 2022, a period that includes the two most severe U.S. surges: the Delta and Omicron waves. The Delta evaluation period spans 15 weeks (Jun 21 - Sep 27, 2021), capturing the rapid rise in late July-Early August, the peak in late August-early September, and the decline throughout September. The second evaluation period spans 10 weeks, primarily focusing on the Omicron wave in the U.S., encompassing the substantial surge in cases in late December 2021, the peak in mid-January 2022, and the subsequent decline in late January 2022. These periods were chosen for model evaluation to test generalizability, as it covered the latest major wave of cases and hospitalizations in the U.S., which challenged forecast models given the explosive spread of the Delta and Omicron variant (Lopez et al. 2024) and the abrupt changes in hospitalization dynamics between the two waves. Although the Pearson correlation between cases and hospitalizations remained strong at 0.74 during the Omicron wave (December 6, 2021, to February 7, 2022)—even higher than during the Delta wave (0.62)—the case-to-hospitalization ratio nearly doubled (28.52 vs. 13.40). This underscores that, while overall trends in cases and hospitalizations can remain broadly aligned, the magnitude shifted significantly during Omicron, with far more infections per resulting hospitalization—a change likely driven by variant-specific factors and the rise of at-home testing. Consequently, forecasting models must focus on direct hospitalization indicators rather than solely relying on infection trends (Fuss et al. 2022).
We evaluate and compare our model performance using three commonly used point-forecast metrics and one probabilistic evaluation score.
@@
Mean Absolute Percentage Error (MAPE) measures accuracy as a percentage and is scale-independent and easy to interpret. However, MAPE may produce infinite and undefined values for zero or close-to-zero true values. As our target is state-level values, it is very unlikely that there will be zero or close-to-zero values. Formally, MAPE at time $t$ is defined as:
@@
In addition to point-based errors, we evaluate the Weighted Interval Score (WIS), a proper scoring rule designed to evaluate probabilistic forecasts that approximates the continuous ranked probability score. This score was adopted by the Hub as the main evaluation metric (Cramer et al. 2022b). WIS rewards forecasts that are both sharp (narrow intervals) and well-calibrated (intervals that cover the truth at multiple quantile levels). Let $m$ be the predictive median and $\{(l_{k},u_{k})\}_{k=1}^{K}$ the $K$ central prediction intervals with nominal coverages $1-\alpha_{k}$. Then
@@ 4.1 Predictive Power of Spatial Features
To evaluate the predictive performance of the Facebook-derived spatiotemporal feature (SPH), we trained three identical multi-horizon ensemble parallel LSTMs: one with all input features (smoothed cases, smoothed hospitalizations, and SPH), one excluding SPH, and a third model incorporating both SPH and SPC. The results in Table 4.1, Fig. 5, and Fig. 10 demonstrate that the model with SPH has less error than the model without SPH across almost the entire 28-day horizon, except the first day in terms of MAE and MAPE, and the first two days in RMSE (The full daily comparison is available in Table C). On average, the spatial model outperforms the non-spatial model by 363 hospitalizations per state during the 28-day forecasting horizon, with the largest difference observed on the $28^{th}$ day, where the model with SPH is approximately 23 hospitalizations per state more accurate in MAE.
Interestingly, the model incorporating both SPH and SPC does not show consistent improvement and, in fact, experiences degradation in performance in most cases, particularly beyond the first four weeks of the evaluation period. During the initial weeks, the SPC-enhanced model demonstrates comparable or slightly better performance, especially until December 27, 2021. However, its performance declines significantly at the turning point of January 7, 2022, and in the subsequent weeks of the decreasing hospitalization trend. This result aligns with the hypothesis that the relationship between infections and hospitalizations shifted significantly during the Omicron wave. The lack of improvement when including SPC suggests that the spatial information embedded in recorded cases during this period does not contribute to the predictive performance of hospitalization models during such dynamic pandemic phases, especially in the presence of at-home testing that is not reported.
@@
Beyond minimizing errors, we also found that the spatial models (with SPH) had a lower variance of 15 ensemble members during the evaluation period. The variances of the spatial model are consistently lower than the non-spatial model, except for the last evaluation week (Table 4.1). The higher variances of the non-spatial model indicate that the model may be overfitting to the noise present in the data. On the contrary, the spatial model is able to capture the trends in data and has better generalizability.
@@ 4.2 Ablation Study of Multi-horizon Ensemble Strategy
To assess the effectiveness of the multi-horizon ensemble strategy, we compared two different forecasting models: the SLSTM (Multi-horizon Ensemble) and the SLSTM (28-Day Direct). The multi-horizon ensemble model leverages an ensemble strategy as described in Section 3.4, where different members of the total 15 members learn on overlapping 7-day, 14-day, and 28-day periods. Conversely, the 28-day direct model consolidates the forecast into a single 28-day output span, with 15 models each predicting the entire 28-day period directly.
The comparative performance of these models is visually presented in Fig. 6. The SLSTM (Multi-horizon Ensemble) consistently outperforms the SLSTM (28-Day Direct) across all time steps. Specifically, the multi-horizon strategy improves over the 28-day direct approach by an average of 167 hospitalizations per state throughout the evaluation period. This is also supported by lower Mean Absolute Percentage Error (MAPE) and Root Mean Squared Error (RMSE) values for the multi-horizon ensemble across the 28-day forecasting period in Fig. 11.
@@
The superior performance of the SLSTM (Multi-horizon Ensemble) can be attributed to its ability to reduce error over segments, as well as mitigate error propagation and drift inherent in longer-term direct forecasts. This strategy ensures a balance between errors at individual time steps and overall consistency for complex forecasting tasks such as predicting hospitalizations over an extended period.
@@ 4.3 Comparison Against Baselines
The results in Table 4.3, and visualizations (in Fig. 7 and Fig. 12) show that SLSTM consistently outperforms all other comparison models across the Delta and Omicron on all four evaluation metrics. Daily comparisons (Table C) for the Delta wave reveal that SLSTM’s advantage is larger for the first two-week horizon, while the performance in week 3 and 4 is only marginally better than that of the COVIDhub-4_week_ensemble. This likely reflects the impact of limited training data for long-lead predictions during pandemic surges on deep learning-based models. During the Omicron wave, SLSTM not only preserved its short‑term advantage but also achieved substantial gains at longer horizons (weeks 3–4) (shown in Fig. C). Specifically, SLSTM is, on average, 15, 30, 47, and 64 hospitalizations per state more accurate than the next best model, COVIDhub-4_week_ensemble, for the $2^{nd},9^{th},16^{th}$, and $23^{rd}$ day, respectively. The COVIDhub-4_week_ensemble has the second smallest error, and COVIDhub-trained_ensemble performs worse than the COVIDhub-baseline in certain periods according to MAE and RMSE. This shows the value of deep learning-based models when sufficient, high-quality training data are available.
@@
Looking at individual forecast dates in the evaluation period, SLSTM outperforms all other models in 7 of 10 forecast dates. The three dates that SLSTM has higher average prediction errors than COVIDhub-4_week_ensemble occur during the increase (December 13, 2021 and December 20, 2021) and the decline (February 7, 2022) in hospitalizations. It is worth noting that SLSTM is substantially more accurate than the other three models around the peak of incident hospitalization during the Omicron wave (the middle of our evaluation period). This is due to comparison models overpredicting the peak, while SLSTM successfully forecasts the beginning of the decline of the Omicron wave.
Beyond point forecasts, SLSTM also excels in probabilistic forecasting. It has the lowest overall WIS in both waves (See Table 4.3). The improvement can reach 37% over the next best model, COVIDhub-4_week_ensemble, in the Omicron wave. When decomposing WIS into three components (shown in Table 4.3, C, and C), SLSTM achieves the smallest dispersion in both waves, indicating tighter and more reliable prediction intervals. Consistent with the point‑based metrics, SLSTM leads on short‑term horizons and maintains a slight advantage over the equally weighted ensemble (COVIDhub‑4_week_ensemble) across the 28‑day window. During the Omicron wave, its under-prediction performance is similar to COVIDhub-4_week_ensemble, but its over-prediction errors are significantly lower, showing fewer high-end misses and better performance in predicting turning point and decreasing phases.
@@
Fig. 8 presents the spatial distribution of prediction error on $2^{nd},9^{th},16^{th}$, and $23^{rd}$ days ahead as forecasted on January 3, 2022. All choropleth maps are normalized by population. Overall, SLSTM shows no systematic bias and has uniformly lower prediction errors across low-, median-, and high-incidence states than COVIDhub-4_week_ensemble and COVIDhub-trained_ensemble on all four dates. The two ensemble models tend to overpredict in states with larger populations, while SLSTM has relatively smaller prediction errors. The patterns are most obvious in the forecasts of the last row ($23^{rd}$ day), with less spatial pattern left in the residuals.
Fig. 9 compares hospitalization forecasts made on January 3, 2022, by SLSTM and the three comparison models. We chose the six states based on their population and hospitalization numbers. While it is evident from Fig. 9 that all models, including SLSTM, tended to overestimate hospitalization at the peak of the Omicron wave, our model was notably able to predict the non-linear changing trend of hospitalizations—transitioning from increasing to decreasing phases. The hospitalization curves of the six states at this time show heterogeneity across the space, as eastern states were already declining, while California and Texas were about to reach their respective peaks.
@@ 4.4 Prediction Intervals and Model Uncertainty
The blue shaded band in Fig. 9 represents the 95% prediction interval of SLSTM, generated using quantiles 0.025 and 0.975 in the model output for each output time-step. The uncertainty of forecasts increases further with the forecast horizon due to the inherent challenges of longer-term predictions. This increase in uncertainty is primarily driven by error propagation as the model increasingly relies on projections rather than direct observations. Additionally, changes in the characteristics of spatial units and variations in interventions over time introduce heterogeneity into the (historical) training and (future) testing sets, further contributing to uncertainty. While this is a perennial challenge in time-series forecasting, incorporating more dynamic features has the potential to help mitigate this effect. Future research should focus on developing more strategies to further reduce forecast uncertainty. Nevertheless, during the evaluation period and for all states, the 95% prediction intervals of SLSTM forecasts include 90.20% of ground truth values, while only 68.63% and 70.24% of ground truth values fall into the 95% prediction intervals of COVIDhub-4_week_ensemble, and COVIDhub-trained_ensemble, respectively. These qualitative findings align with the quantitative WIS results reported in Table 4.3: SLSTM achieves the smallest dispersion and over-prediction penalties, confirming that its prediction intervals are tighter and more reliable. Together, these metrics demonstrate SLSTM’s superior ability to quantify the prediction uncertainty and capture true values in the prediction intervals.
@@ 5 Discussion
Our evaluation results highlight the novel approach of incorporating spatiotemporal features derived from Facebook’s Social Connectedness Index into a multi-horizon, parallel LSTM framework. This unique integration significantly improves predictive performance and reduces variance and uncertainty in the hospitalization projection, particularly over longer-term horizons.
@@
Human mobility and the resulting interactions contribute to the transmission dynamics of infectious diseases (Changruenngam et al. 2020). From a geospatial modeling point of view, this is similar to the spatial spillover effect in spatial lag modeling, where the target variable also influences the target variable in connected spatial units. However, accessing human mobility data can be expensive and limited due to privacy concerns and the cost of data, and it is simply impossible for many places and countries. At the time of this writing, many mobility datasets that temporarily opened up during the pandemic are unavailable even for the US post-pandemic. By contrast, Facebook’s SCI is released and updated annually, remains freely accessible, and covers county-, state-, and country-level in more than 200 countries and territories, providing a stable, no-cost proxy for human interaction that persists beyond the pandemic period (Bailey et al. 2017). Beyond infectious disease modeling, SCI has been widely applied in studies of COVID-19 vaccine uptake (Basu et al. 2025), residential segregation (Nilforoshan et al. 2023), international migration (Minora et al. 2022), and eco-innovation partnerships (Basak et al. 2024), highlighting its broad applicability across public health, socio-economic, and behavioral domains.
@@
Human mobility from spatial unit to spatial unit is highly correlated with social networks, and both follow spatial constraints meaning that the majority of our trips or communications happen locally, with occasional longer journeys or calls to friends residing at a distance (Deville et al. 2016). Though the Social Connectedness Index (SCI) is a relatively static measurement based on social media friendships among pairs of spatial units, it is a good alternative proxy for mobility data such as those distributed by SafeGraph (Ilin et al. 2021)—an advantage that is critical for low- and middle-income regions where commercial mobility data are either cost-prohibitive or entirely absent. In previous research, SCI-derived features demonstrated slightly stronger predictive power in short-term forecasting than SafeGraph’s mobility data, albeit using tree-based ensemble models (Vahedi et al. 2021).
@@
One of the advantages of the statistical forecasting of infectious diseases, such as our proposed model, is that such approaches can capture spatial heterogeneity and patterns in the data, even when variables are not directly included as predictor features. For instance, although vaccination clearly mitigates Omicron severity, reliable state-level vaccination data were unavailable during our study window (July 2020 - March 2022). Many jurisdictions did not publish daily dose counts or demographic breakdowns, and CDC’s data aggregation pipeline documented frequent gaps and de-duplication challenges (Bradley et al. 2021, Scharf et al. 2024). Therefore, rather than injecting a noisy, sparsely reported vaccination variable, our data-driven approach does not include predictor variables such as vaccination rates, especially since the vaccine versions and the length and extent of immunity are highly variable. Instead, we rely on incident cases, hospitalizations, and a SCI-derived spatiotemporal hospitalization lag feature, SPH. Because higher vaccine coverage lowers the infection-to-hospitalization ratio, those declined hospitalization signals are encoded in the time series and then propagated to socially connected states through the SCI weights, allowing SPH to capture vaccine-driven heterogeneity implicitly (Alvarez-Zuzek et al. 2022, Aslim et al. 2024). Emerging evidence shows that county‑level vaccination uptake is strongly shaped by friendship networks, implying that SCI already embeds vaccination dynamics implicitly (Basu et al. 2025). Similarly, quantifying non-pharmaceutical interventions (NPIs) such as quarantine compliance, social distancing, and changes in health care provision is difficult and, therefore, challenging to incorporate into models directly (Iezadi et al. 2021). Instead, the statistical model learns from the patterns in data, where NPI efficacy is reflected in the time-series inputs of cases and hospitalizations in spatial units.
@@
Compared to the compartmental models (i.e., the SIR family), the statistical approaches of deep learning update complex models with many learnable parameters with non-linear activation functions by feeding training data and backpropagation instead of specifying differential equations and calibrating limited parameters. This flexibility means the neural network can internalize time-varying factors such as emerging variants, shifts in mobility, and spatially clustered vaccine uptake via SPH, without explicit corresponding covariates. Due to the high variability of SARS-CoV-2 variants and heterogeneity across space and time, the key indicator, $R_{0}$, varies between variants (Du et al. 2022), spatial units (Thomas et al. 2020), and age groups (Monod et al. 2021). This is another indication that compartmental models are not complex enough to capture the heterogeneity of the population as well as virus variants across space and time, compared to deep learning models. However, $R_{0}$ is essential for tracking changes in disease transmissibility, allowing decision-makers to assess the effectiveness of interventions. Other parameters make it possible to simulate different scenarios for policy-making, such as the slow waning of immunity and absence of a new variant, and fast waning, and with the emergence of a new variant (Borchering 2021).

### v2

@@ 3.5 Model Training
We employed a walk-forward validation approach to train the model, utilizing all available data up to and including the forecast date, which is made accessible before the Hub submission deadline every week. Models for each forecast date were trained using the most recent 15 months of data, capturing the latest transmission patterns, hospitalization dynamics, and population immunity. We used spatial cross-validation for training our models, and the final testing was conducted purely on forecasting dates during periods that occurred after the training set period. We configured our training strategy such that for each forecast date, it selected three states exhibiting the lowest, median, and highest hospitalization rates (a criterion that changes weekly, so different states occupy each role over time) to serve as our validation for early stopping. This spatial validation strategy ensures that the model is trained while consistently validated against a diverse range of epidemic dynamics- specifically, states at increasing, decreasing, and persisting stages of epidemic spread (Lever et al. 2016). This ensures that the model is not overfitting to increasing or decreasing trends. The remaining states, with varying hospitalization rate trends are used for training the model. This method allows us to implement early stopping to prevent overfitting and test the model’s generalizability across a representative cross-section of epidemic conditions. After training is completed, the model is tested on time periods following the training period (with no overlap), to ensure the model’s ability to forecast the future.
During training, the time-series SPH values, along with incident cases and hospitalization rates, were first normalized to the range [-1, 1] using the MinMax Scaling technique. These normalized features were then sequentially fed into the LSTM model, which processes the three input features over successive time steps. This integration enhances the model’s ability to predict hospitalizations based not only on historical patterns within the same state, but also on historical patterns across connected states, with the strength of connectedness measured by social connectivity.
@@ 3.6 Model Evaluation
While SLSTM was employed during various phases of the pandemic to submit forecasts to the Hub, for the results presented in this paper, we utilized data from July 27, 2020, to March 7, 2022, a period that includes the two most severe U.S. surges: the Delta and Omicron waves. The Delta evaluation period spans 15 weeks (Jun 21 - Sep 27, 2021), capturing the rapid rise from late July to early August, the peak from late August to early September, and the decline throughout September. The second evaluation period spans 10 weeks, primarily focusing on the Omicron wave in the U.S., encompassing the substantial surge in cases in late December 2021, the peak in mid-January 2022, and the subsequent decline in late January 2022. These periods were chosen for model evaluation to test generalizability, as it covered the latest major wave of cases and hospitalizations in the U.S., which challenged forecast models given the explosive spread of the Delta and Omicron variant (Lopez et al. 2024) and the abrupt changes in hospitalization dynamics between the two waves. Although the Pearson correlation between cases and hospitalizations remained strong at 0.74 during the Omicron wave (December 6, 2021, to February 7, 2022)—even higher than during the Delta wave (0.62)—the case-to-hospitalization ratio nearly doubled (28.52 vs. 13.40). This underscores that, while overall trends in cases and hospitalizations can remain broadly aligned, the magnitude shifted significantly during Omicron, with far more infections per resulting hospitalization—a change likely driven by variant-specific factors and the rise of at-home testing. Consequently, forecasting models must focus on direct hospitalization indicators rather than solely rely on infection trends (Fuss et al. 2022).
We evaluated and compared our model performance using three commonly used point-forecast metrics and one probabilistic evaluation score.
@@
Mean Absolute Percentage Error (MAPE) measures accuracy as a percentage and is scale-independent and easy to interpret. However, MAPE may produce infinite and undefined values for zero or close-to-zero true values. As our target is state-level values, it is highly unlikely that there will be zero or close-to-zero values. Formally, MAPE at time $t$ is defined as:
@@
In addition to point-based errors, we evaluated the Weighted Interval Score (WIS), a proper scoring rule designed to evaluate probabilistic forecasts that approximates the continuous ranked probability score. This score was adopted by the Hub as the main evaluation metric (Cramer et al. 2022b). WIS rewards forecasts that are both sharp (narrow intervals) and well-calibrated (intervals that cover the truth at multiple quantile levels). Let $m$ be the predictive median and $\{(l_{k},u_{k})\}_{k=1}^{K}$ the $K$ central prediction intervals with nominal coverages $1-\alpha_{k}$. Then
@@ 4.1 Predictive Power of Spatial Features
To evaluate the predictive performance of the Facebook-derived spatiotemporal feature (SPH), we trained three identical multi-horizon ensemble parallel LSTMs: one with all input features (smoothed cases, smoothed hospitalizations, and SPH), one excluding SPH, and a third model incorporating both SPH and SPC. The results in Table 4.1, Fig. 5, and Fig. 10 demonstrate that the model with SPH has lower errors than the model without SPH across almost the entire 28-day horizon, except the first day in terms of MAE and MAPE, and the first two days in RMSE (the full daily comparison is available in Table C). On average, the spatial model outperformed the non-spatial model by 363 hospitalizations per state over the 28-day forecasting horizon, with the largest difference observed on the $28^{th}$ day, where the model with SPH was approximately 23 hospitalizations per state more accurate in MAE.
Interestingly, the model incorporating both SPH and SPC did not show consistent improvement and, in fact, suffered degradation in performance in most cases, particularly beyond the first four weeks of the evaluation period. During the initial weeks, the SPC-enhanced model demonstrated comparable or slightly better performance, especially until December 27, 2021. However, its performance declined significantly at the turning point of January 7, 2022, and in the subsequent weeks of the decreasing hospitalization trend. This result aligns with the hypothesis that the relationship between infections and hospitalizations shifted markedly during the Omicron wave. The lack of improvement when including SPC suggests that the spatial information embedded in recorded cases during this period did not contribute to the predictive performance of hospitalization models during such dynamic pandemic phases, especially in the presence of at-home testing.
@@
Beyond minimizing errors, we also found that the spatial models (with SPH) had a lower variance across the 15 ensemble members during the evaluation period. The variances of the spatial model were consistently lower than those of the non-spatial model, except for the final evaluation week (Table 4.1). The higher variances of the non-spatial model indicate that the model may be overfitting to the noise present in the data. On the contrary, the spatial model was able to capture the trends in data and has better generalizability.
@@ 4.2 Ablation Study of Multi-horizon Ensemble Strategy
To assess the effectiveness of the multi-horizon ensemble strategy, we compared two different forecasting models: the SLSTM (Multi-horizon Ensemble) and the SLSTM (28-Day Direct). The multi-horizon ensemble model leveraged an ensemble strategy as described in Section 3.4, where different members of the total 15 members learn on overlapping 7-day, 14-day, and 28-day periods. Conversely, the 28-day direct model consolidated the forecast into a single 28-day output span, with 15 models each predicting the entire 28-day period directly.
The comparative performance of these models is visually presented in Fig. 6. The SLSTM (Multi-horizon Ensemble) consistently outperformed the SLSTM (28-Day Direct) across all time steps. Specifically, the multi-horizon strategy improved over the 28-day direct approach by an average of 167 hospitalizations per state throughout the evaluation period. This was also supported by lower Mean Absolute Percentage Error (MAPE) and Root Mean Squared Error (RMSE) values for the multi-horizon ensemble across the 28-day forecasting period in Fig. 11.
@@
The superior performance of the SLSTM (Multi-horizon Ensemble) can be attributed to its ability to reduce error over segments, as well as mitigate error propagation and drift inherent in longer-term direct forecasts. This strategy ensures a balance between errors in individual time steps and overall consistency for complex forecasting tasks such as predicting hospitalizations over an extended period.
@@ 4.3 Comparison Against Baselines
The results in Table 4.3 and visualizations (in Fig. 7 and Fig. 12) show that SLSTM consistently outperformed all other comparison models across the Delta and Omicron in all four evaluation metrics. Daily comparisons (Table C) for the Delta wave reveal that SLSTM’s advantage was larger for the first two-week horizon, while the performance in week 3 and 4 was only marginally better than that of the COVIDhub-4_week_ensemble. This likely reflects the impact of limited training data for long-lead predictions during pandemic surges on deep learning-based models. During the Omicron wave, SLSTM not only preserved its short‑term advantage but also achieved substantial gains at longer horizons (weeks 3–4) (shown in Fig. C). Specifically, SLSTM was, on average, 15, 30, 47, and 64 hospitalizations per state more accurate than the next best model, COVIDhub-4_week_ensemble, for the $2^{nd},9^{th},16^{th}$, and $23^{rd}$ day, respectively. The COVIDhub-4_week_ensemble had the second smallest error, and COVIDhub-trained_ensemble performed worse than the COVIDhub-baseline in certain periods according to MAE and RMSE. This shows the value of deep learning-based models when sufficient, high-quality training data are available.
@@
Looking at individual forecast dates within the evaluation period, SLSTM outperformed all other models in 7 of 10 forecast dates. The three dates that SLSTM had higher average prediction errors than COVIDhub-4_week_ensemble occur during the increase (December 13, 2021, and December 20, 2021) and the decline (February 7, 2022) in hospitalizations. It is worth noting that SLSTM was substantially more accurate than the other three models around the peak of incident hospitalization during the Omicron wave (the middle of our evaluation period). This is due to comparison models overpredicting the peak, while SLSTM successfully forecasted the beginning of the decline of the Omicron wave.
Beyond point forecasts, SLSTM also excelled in probabilistic forecasting. It had the lowest overall WIS in both waves (See Table 4.3). The improvement can reach 37% over the next best model, COVIDhub-4_week_ensemble, in the Omicron wave. When decomposing WIS into three components (shown in Table 4.3, C, and C), SLSTM achieved the smallest dispersion in both waves, indicating tighter and more reliable prediction intervals. Consistent with the point‑based metrics, SLSTM led on short‑term horizons and maintained a slight advantage over the equally weighted ensemble (COVIDhub‑4_week_ensemble) across the 28‑day window. During the Omicron wave, its under-prediction performance was similar to COVIDhub-4_week_ensemble, but its over-prediction errors were significantly lower, showing fewer high-end misses and better performance in predicting turning point and decreasing phases.
@@
Fig. 8 presents the spatial distribution of the prediction errors on $2^{nd},9^{th},16^{th}$, and $23^{rd}$ days ahead as forecasted on January 3, 2022. All choropleth maps are normalized by population. Overall, SLSTM showed no systematic bias and had uniformly lower prediction errors across low-, median-, and high-incidence states than COVIDhub-4_week_ensemble and COVIDhub-trained_ensemble on all four dates. The two ensemble models tended to overpredict in states with larger populations, while SLSTM had relatively smaller prediction errors. The patterns were most obvious in the forecasts for the last row ($23^{rd}$ day), with less spatial pattern left in the residuals.
Fig. 9 compares hospitalization forecasts made on January 3, 2022, by SLSTM and the three comparison models. We chose the six states based on their population and hospitalization numbers. While it is evident from Fig. 9 that all models, including SLSTM, tended to overestimate hospitalization at the peak of the Omicron wave, our model was notably able to predict the non-linear changing trend of hospitalizations—transitioning from increasing to decreasing phases. The hospitalization curves of the six states at this time showed heterogeneity across the space, as eastern states were already declining, while California and Texas were about to reach their respective peaks.
@@ 4.4 Prediction Intervals and Model Uncertainty
The blue shaded band in Fig. 9 represents the 95% prediction interval of SLSTM, generated using quantiles 0.025 and 0.975 in the model output for each output time-step. The uncertainty of forecasts increased further with the forecast horizon due to the inherent challenges of longer-term predictions. This increase in uncertainty is primarily driven by error propagation as the model increasingly relies on projections rather than direct observations. Additionally, changes in the characteristics of spatial units and variations in interventions over time introduced heterogeneity into the (historical) training and (future) testing sets, further contributing to uncertainty. While this is a perennial challenge in time-series forecasting, incorporating more dynamic features has the potential to help mitigate this effect. Future research should focus on developing more strategies to further reduce forecast uncertainty. Nevertheless, during the evaluation period and for all states, the 95% prediction intervals of SLSTM forecasts included 90.20% of ground truth values, while only 68.63% and 70.24% of ground truth values fell into the 95% prediction intervals of COVIDhub-4_week_ensemble, and COVIDhub-trained_ensemble, respectively. These qualitative findings align with the quantitative WIS results reported in Table 4.3: SLSTM achieved the smallest dispersion and over-prediction penalties, confirming that its prediction intervals are tighter and more reliable. Together, these metrics demonstrated SLSTM’s superior ability to quantify the prediction uncertainty and capture true values in the prediction intervals.
@@ 5 Discussion
Our evaluation results highlighted the novel approach of incorporating spatiotemporal features derived from Facebook’s Social Connectedness Index into a multi-horizon, parallel LSTM framework. This unique integration significantly improved predictive performance and reduced variance and uncertainty in the hospitalization projection, particularly over longer-term horizons.
@@
Human mobility and the resulting interactions contribute to the transmission dynamics of infectious diseases (Changruenngam et al. 2020). From a geospatial modeling point of view, this is similar to the spatial spillover effect in spatial lag modeling, where the target variable also influences the target variable in connected spatial units. However, accessing human mobility data can be expensive and limited due to privacy concerns and the high cost of data, making it simply impossible for many places and countries. At the time of this writing, many mobility datasets that were temporarily made available during the pandemic are still unavailable, even in the US post-pandemic. In contrast, Facebook’s SCI is released and updated annually, remains freely accessible, and covers county-, state-, and country-level in more than 200 countries and territories, providing a stable, no-cost proxy for human interaction that persists beyond the pandemic period (Bailey et al. 2017). Beyond infectious disease modeling, SCI has been widely applied in studies of COVID-19 vaccine uptake (Basu et al. 2025), residential segregation (Nilforoshan et al. 2023), international migration (Minora et al. 2022), and eco-innovation partnerships (Basak et al. 2024), highlighting its broad applicability across public health, socio-economic, and behavioral domains.
@@
Human mobility from one spatial unit to another is highly correlated with social networks, and both follow spatial constraints, meaning that the majority of our trips or communications occur locally, with occasional longer journeys or calls to friends residing at a distance (Deville et al. 2016). Though the Social Connectedness Index (SCI) is a relatively static measurement based on social media friendships among pairs of spatial units, it is a good alternative proxy for mobility data such as those distributed by SafeGraph (Ilin et al. 2021)—an advantage that is critical for low- and middle-income regions where commercial mobility data are either cost-prohibitive or entirely absent. In previous research, SCI-derived features demonstrated slightly stronger predictive power in short-term forecasting than SafeGraph’s mobility data, albeit using tree-based ensemble models (Vahedi et al. 2021).
@@
One of the advantages of statistical forecasting for infectious diseases, such as our proposed model, is that such approaches can capture spatial heterogeneity and patterns in the data, even when variables are not directly included as predictor features. For instance, although vaccination clearly mitigates Omicron severity, reliable state-level vaccination data were unavailable during our study window (July 2020 - March 2022). Many jurisdictions did not publish daily dose counts or demographic breakdowns, and the CDC’s data aggregation pipeline documented frequent gaps and de-duplication challenges (Bradley et al. 2021, Scharf et al. 2024). Therefore, rather than injecting a noisy, sparsely reported vaccination variable, our data-driven approach did not include predictor variables such as vaccination rates, especially since the vaccine versions, as well as the length and extent of immunity, are highly variable. Instead, we relied on incident cases, hospitalizations, and a SCI-derived spatiotemporal hospitalization lag feature, SPH. Because higher vaccine coverage lowers the infection-to-hospitalization ratio, those declined hospitalization signals are encoded in the time series and then propagated to socially connected states through the SCI weights, allowing SPH to capture vaccine-driven heterogeneity implicitly (Alvarez-Zuzek et al. 2022, Aslim et al. 2024). Emerging evidence shows that county‑level vaccination uptake is strongly shaped by friendship networks, implying that SCI already embeds vaccination dynamics implicitly (Basu et al. 2025). Similarly, quantifying non-pharmaceutical interventions (NPIs) such as quarantine compliance, social distancing, and changes in health care provision is difficult and, therefore, challenging to incorporate into models directly (Iezadi et al. 2021). Instead, the statistical model learns from the patterns in data, where NPI efficacy is reflected in the time-series inputs of cases and hospitalizations in spatial units.
@@
Compared to compartmental models (i.e., the SIR family), the statistical approaches of deep learning update complex models with many learnable parameters using non-linear activation functions by feeding training data and employing backpropagation, rather than specifying differential equations and calibrating limited parameters. This flexibility enables the neural network to internalize time-varying factors, such as emerging variants, shifts in mobility, and spatially clustered vaccine uptake, via SPH, without requiring explicit corresponding covariates. Due to the high variability of SARS-CoV-2 variants and heterogeneity across space and time, the key indicator, $R_{0}$, varies between variants (Du et al. 2022), spatial units (Thomas et al. 2020), and age groups (Monod et al. 2021). This is another indication that compartmental models are not complex enough to capture the heterogeneity of the population as well as virus variants across space and time, compared to deep learning models. However, $R_{0}$ is crucial for tracking changes in disease transmissibility, enabling decision-makers to evaluate the effectiveness of interventions. Other parameters make it possible to simulate different scenarios for policy-making, such as the slow waning of immunity and absence of a new variant, and fast waning, and with the emergence of a new variant (Borchering 2021).